OpenStudy (anonymous):

Rewrite in vertex form: y = x^2 + 14x +50

OpenStudy (anonymous):

First thing you must do is complete the square.

OpenStudy (anonymous):

\[ (x+a)^2 = x^2+2ax + a^2 \]

OpenStudy (anonymous):

Ugh people keep saying that bu I dont know what that means

OpenStudy (anonymous):

Basically in general:\[ (x+a)^2 = x^2+\color{red}{2a}x + a^2 \]In our case we have: \[ (x+a)^2 = x^2+\color{red}{14}x + 50 \]This means that \(2a=14\), which means that \(a=7\) and \(a^2=49\).

OpenStudy (anonymous):

So we are close to being a square, but not quite.

OpenStudy (anonymous):

\[ (x+7)^2 = x^2+14+49 \]This means that \[ x^2+14x+50 = (x+7)^2+1 \]This is what is meant by completing the square.

OpenStudy (anonymous):

Anyway: \[ y = (x+7)^2+1 \]This is the vertex form.

OpenStudy (anonymous):

thank you so much! that was amazing!

OpenStudy (anonymous):

I made a few typos...

OpenStudy (anonymous):

But do you understand how to complete the square?

OpenStudy (anonymous):

Yes. So then thats it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

In general \[ x^2+bx \to (x+b/2)^2-(b/2)^2 \]

OpenStudy (anonymous):

So, using this general rule...\[ x^2+14x = (x+7)^2-49 \]And we can plug it into our previous equation: \[ y=x^2+14x+50 = (x+7)^2-49+50 =(x+7)^2+1 \] This is fasterway to do it, now that you understand it better.